OpenStudy (anonymous):

The annual Gross Domestic Product (GDP) of a country is the value of all of the goods and services produced in the country during a year. During the period 1985-1999, the Gross Domestic Product of the United States grew about 3.2% per year, measured in 1996 dollars. In 1985, the GDP was $577 billion. In what year will the GDP reach $2 trillion? You just need to give the year, not part of a year

OpenStudy (helder_edwin):

if the GDP is growing at 3.2% a year then \[ \large GDP_{n+1}=GDP_n+3.2\%GDP_n=(1+3.2\%)GDP_n \]

OpenStudy (anonymous):

omg

OpenStudy (helder_edwin):

u are studing economics this is simple arithmetic

OpenStudy (helder_edwin):

say u have 100$ and your money grows at 3% every week after a week how much do u have?

OpenStudy (anonymous):

this is pre-cal it uses the y function formula

OpenStudy (helder_edwin):

@brendas

OpenStudy (helder_edwin):

@brendas are u gonna take part of this?

OpenStudy (anonymous):

sorry im back

OpenStudy (anonymous):

what does the n stand for

OpenStudy (helder_edwin):

the year

OpenStudy (anonymous):

ok hold on let me see if i can work it out

OpenStudy (anonymous):

i cant work it

OpenStudy (helder_edwin):

say u have 100$ and your money grows at 3% every week after a week how much do u have?

OpenStudy (anonymous):

100 * 3= 300 300/100= 3 3+100=103

OpenStudy (helder_edwin):

yes that is what the formula says after one year u have what u had in the beginning plus the interests gained

OpenStudy (helder_edwin):

so in general \[ \large GDP_n=(1+3.2\%)^{n-1985}GDP_{1985} \]

OpenStudy (helder_edwin):

u dont know n but u know that \[ \large GDP_n=\$2\text{ trillion} \]

OpenStudy (helder_edwin):

so the last one becomes an equation for n

OpenStudy (anonymous):

right

OpenStudy (anonymous):

??